OpenStudy (anonymous):

Estimate by rounding to the nearest half. five and four-sevenths plus three and seven-eighths minus four and two-ninths three and one-half four and one-half five and one-half six and one-half

OpenStudy (anonymous):

can you type it in numbers?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

include parenthesis if there is any

OpenStudy (anonymous):

2nd one is first

OpenStudy (anonymous):

k one sec

OpenStudy (anonymous):

4 and a half I think

OpenStudy (anonymous):

D

OpenStudy (anonymous):

The answer is 5 1/2. 39/7 + 31/8 - 38/9 = 5.22

OpenStudy (anonymous):

@kadonagaw your supposed to find the common denominator first. can't just subtract...

OpenStudy (anonymous):

so which one is it? >.>

OpenStudy (anonymous):

4 and a half is what I think. But if you want to make sure wait till yours gets bumped up

OpenStudy (anonymous):

@e.mccormick

OpenStudy (anonymous):

five and four-sevenths plus three and seven-eighths minus four and two-ninths \[\huge 5 \frac{4}{7} + 3\frac{7}{8} - 4\frac{2}{9}\] Start by getting them into improper fraction form

OpenStudy (anonymous):

39 + 31 - 38

OpenStudy (anonymous):

32

OpenStudy (anonymous):

you can't just add fractions like that. 5 and 4/7 becomes 39/7

OpenStudy (anonymous):

-.-

OpenStudy (anonymous):

@zepdrix \

OpenStudy (anonymous):

This is basic math. Convert them to improper fractions. At least put some effort in ~_~

OpenStudy (anonymous):

well i need an answer asap. im on a timed test i can only take once